• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 15:58:20 on July16,2003

    Entry number 104068

    keyword=shift summary

    We were ramping up the Septum (goal=198A) at the beginning of the shift;

    From 8:20 to 10:00 target temperature test, results were analyzed by Vince, see halog #104065;

    9:00am controlled access, Bogdan and Rob went down to check the R-arm gas cherenkov cable ADC #3; Later we used cosmic run 2326 to check gas cherenkov, ADC#3 looks fine, see halog #104044;

    10:00, Septum was ready. R-arm Q2 tripped two times since beginning of the shift, both were reset from upstairs;

    11:00-12:40, Jian-Ping and Xiaochao went down to the hall to move the downstream carbon foil. For details please see J.P.'s halog entry #104064. Then MCC need to check something in the Hall which caused a small delay in the beam delivery;

    1:30, Hall is back to beam permit. But they need to cycle the transport dipole magnet (? might also be a communication problem because I forgot to tell them that we have finished the target movement);

    2:05pm, 2uA to Hall A. But beam position was BPMA: x=-1.0, y=0.0 and BPMB: x=-1.2, y=-2.0. This means the beam has an angle in vertical position and will (probablly) hit the target ladder (again). The raster is on as well (usually should be off). I called MCC and asked them to adjust BPMB y to 0.0. Meantime, we noticed that the Septum tripped again. Looks like there is a correlation between Septum trip and beam position (angle in vertical position). See halog #104062 by Vince.

    3:00pm, controlled access again, various people went down, Jian-Ping, Alexandre (installing BeO target); Rob (investigating L-arm S2), Ed (R-arm Q2); Jack Segal (checking ADCs on the beamline), and maybe more;

    4:00pm, still controlled access...